OpenStudy (anonymous):

given the function y=(xcubed -1)*(1+x)^1/2 then dy/dx is?

OpenStudy (mimi_x3):

is it \[ \large y = (x^{2} -1)^{ \frac{1}{2} } * (1+x)^{1/2} \]??

OpenStudy (diyadiya):

\[\large y = (x^{3} -1) * (1+x)^{1/2} ~~~ ?\]

OpenStudy (mimi_x3):

woops sorry! i read it wrong

OpenStudy (anonymous):

y=root(x+1)(x-1)(x+1) y=(x+1)*root(x-1) dy/dx=root(x-1)*1+(x+1)*d/dx of root(x-1) dy/dx=root(x-1)*1+(x+1)*i/root(x-1)

OpenStudy (mimi_x3):

well if that is the function; to differentiate it you can use the product rule u = x^3-1 u' = 3x^2 v = (1+x)^(1/2) v' = 1/2(1+x)^(-1/2) Are you able to proceed?